OpenStudy (anonymous):

use a truth table to represent the statement, "if x^2<100, then x<10." interpret the table for this statement.

OpenStudy (anonymous):

it should actually read: \[x^2 \le 100 , x \le 10 \]

OpenStudy (anonymous):

that rather than x^2<100 and x<10

OpenStudy (anonymous):

whats ur question?

OpenStudy (anonymous):

use a truth table to represent the statement, "if x^2<100, then x<10." interpret the table for this statement.

OpenStudy (anonymous):

that is the question

OpenStudy (anonymous):

If \(p \to q\) ? Let p=x^2<100 Let q = x<10

OpenStudy (anonymous):